OpenStudy (anonymous):

Without dividing, do you expect 5/8÷11/16 to be greater than or less than 1? Explain.

OpenStudy (anonymous):

This is pretty easy....Think about what happens if you divide by a number bigger than the first number....say 5 ÷ 10

OpenStudy (anonymous):

You'll always get a value less than one

OpenStudy (anonymous):

its less than 1 . 5/8= 10/16 10/16 divided by 11/16 is <1

OpenStudy (anonymous):

5/8= 10/16 and 11 /16 is less than 1 because it is 10/11

OpenStudy (anonymous):

sunnyd48, we know

OpenStudy (anonymous):

hehe